# SC-Diff: 3D Shape Completion with Latent Diffusion Models

### v1

@@ 5 Conclusion
In this work, we have developed a latent diffusion-based approach for 3D shape completion, integrating a latent diffusion model with a VQ-VAE. When enhanced with 2D losses, this combination effectively compresses TSDF voxel grids into a compact latent space. In that way, our approach surpasses typical resolution limits seen in current diffusion-based methods, achieving $64^{3}$ resolution and generalizing across multiple object classes, in contrast to class-specific models like DiffComplete [8]. By incorporating a spatially consistent conditioning mechanism we can generate diverse and realistic shapes from partial scans, providing a significant advantage over other state-of-the-art methods such as SDFusion [6] and DiffRF [23], which tackle the shape completion task by filling out artificially cropped out boxes.
Despite these advancements, challenges remain, particularly in processing complex scenes at higher resolutions. Future work will focus on addressing these limitations. Integrating preprocessing steps for accurate object bounding box and pose determination, as seen in existing methods  [24, 16, 26], could expand the method’s usability in practical scenarios beyond the positioning of single objects in a canonical space. Additionally, while diffusion models offer high-quality sample generation, their relatively slow sampling speed is a notable drawback. Emerging research in areas like non-Markovian diffusion processes  [34], progressive distillation  [33], and consistency models  [35] present potential pathways to improve efficiency in conditional generation across various data modalities, including 3D shapes.
@@
Appendix 0.A Detailed Network Architecture
@@
Our diffusion model is fundamentally a 3D version of a ControlNet [42]. Similar to DiffComplete [8] and SDFusion [6], our neural network is implemented using a 3D U-Net [32] with an encoder, a middle block, and a decoder, which is connected to the encoder through skip connections. Both the encoder and decoder contain ResNetBlocks. In the case of the encoder, these are accompanied by downsampling blocks, while in the decoder, they are accompanied by up-sampling convolution layers. The encoder, decoder, and middle block include two spatial transformers each. Images are encoded using CLIP encoder [27], and diffusion timesteps are encoded with a 2-layer MLP using positional encoding. Similarly, the control branch shares the same architecture as the diffusion model up to the middle block. Then, it uses 3D convolutional projection layers to feed partial shape’s $\mathbf{c}$ intermediate representation into the diffusion model spatially consistently. (See Fig. 7).
Appendix 0.B Hyperparameter Values
The Tab. 6 shows the hyperparameters chosen in our experiments.
@@ Appendix 0.B Hyperparameter Values
TSDF Truncation threshold $thresh$ $3.0$
@@
Feature size of the VQ-VAE’s latent space $D$ $3$
@@
3D Commitment Loss weight for VQ-VAE $\beta$ $0.5$
2D Reconstruction Loss weight for VQ-VAE $\gamma_{R}$ $0.4$
2D Adversarial Loss weight for VQ-VAE $\gamma_{A}$ $0.4$
@@
Diffusion timesteps at inference (DDIM [34]) $T_{inf}$ $100$
@@
Batch Size $bs$ $8$
Learning Rate VQ-VAE $lr_{vqvae}$ $1e^{-4}$
Learning Rate Diffusion Model $lr_{diff}$ $2.5e^{-5}$
Appendix 0.C Forward and backward diffusion processes
As shown in Fig. 8, the forward diffusion process is a fixed Markov chain that gradually corrupts a latent code $\mathbf{z}_{0}$ by adding Gaussian noise with a given variance schedule. Transitions are defined by $q(\mathbf{z}_{t}|\mathbf{z}_{t-1})$ in Eq. 4. We train a diffusion model to learn the reverse denoising process with transitions $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})$ as shown in Eq. 5
Figure 8: Forward Diffusion Process. After passing them through our pretrained decoder $\mathcal{D}$, we illustrate how the shapes recovered from the corrupted latent codes look.
Appendix 0.D Effect of 2D Loss on VQ-VAE Performance
Figure 9: The introduction of 2D losses improves the VQ-VAE’s reconstruction, thus improving the learning of the latent space.
Appendix 0.E Effect of different conditioning types
Figure 10: Comparison of different types of conditioning.
Appendix 0.F More multimodal results
Fig. 11 shows more multimodal predictions our model performs.
Figure 11: Further multimodal predictions performed by our model
Appendix 0.G Failure Cases
To evaluate the instances where our model does not perform as expected, we carry out five predictions for a given input and select the one with the highest L1 error. This approach enables us to obtain predictions like those shown in Fig. 12. Owing to the intrinsic ambiguity of the shape completion task and the significant incompleteness of some partial scans, the predicted shapes might not only appear unrealistic but also be completed as entirely different objects. This issue partly stems from the fact that our model is class-agnostic.
Figure 12: Examples of instances where our model failed to accurately predict high-fidelity and realistic shapes. a) The model predicts the shape as a type of container, even though the ground-truth shape is a lamp. b) Our latent diffusion model unnecessarily predicts a more complex shape than the ground truth. c) and d) Given the high level of incompleteness of the input, the model completes the shape as a completely different class; for instance, it predicts a table and sofa instead of the actual ground truth, which are a lamp and cabinet

### v2

@@ 5 Conclusion
In this work, we presented a novel framework for 3D shape completion that unifies latent diffusion models with both 2D and 3D conditioning in a spatially consistent manner. By jointly training a VQ-VAE with 2D and 3D supervision, we compress TSDF voxel grids into a compact discrete latent space, enabling $64^{3}$ resolution reconstructions while reducing GPU memory consumption by $30\%$ compared to state-of-the-art approaches. Beyond architectural efficiency, our method advances conditional generation by combining cross-attention with ControlNet-inspired guidance, allowing flexible integration of complementary modalities. To better capture real-world challenges, we introduce a training strategy that simulates realistic partial scans directly from 3D datasets without prior assumptions on input structure. This results in a robust multimodal framework capable of class-agnostic generation while producing diverse and high-fidelity shape completions. Comprehensive experiments on both synthetic and real-world benchmarks confirm that our approach achieves state-of-the-art performance, reducing $l_{1}$ reconstruction error by $12\%$ over class-specific baselines and by $47\%$ over class-agnostic models. These results highlight the advantages of our method in overcoming resolution bottlenecks and improving the realism of 3D completions. Despite these contributions, challenges remain in extending the method to complex scenes. Future work will explore recent advances in efficient diffusion, such as non-Markovian sampling, progressive distillation [37], and consistency models [39]. We believe these avenues will further enhance the practicality of latent diffusion for real-world 3D applications, enabling high-quality, scalable shape completion.
@@
I VQ-VAE Ablations
Figure 7: The introduction of 2D losses produces much cleaner reconstructions in the VQ-VAE, enabling modeling in a compact latent space while preserving crisp geometric details.
II Effect of Different Conditioning Types
Figure 8: Comparison of conditioning modalities. Using only TSDF, the diffusion model successfully completes partial shapes but lacks fine detail. When conditioned on both TSDF and RGB, the model exploits appearance cues to produce sharper and more detailed reconstructions.
III Preprocessing
We employed the ShapeNetCore dataset [6], which contains 55 categories of common objects. For our model, we required data tuples $(\mathbf{X},\mathbf{c},\mathbf{I})$ including complete and partial TSDF grids and RGB images. Given that ShapeNet provides objects’ meshes and textures, we converted these into the necessary TSDF voxel grids. Our process involved rendering 20 views (RGB+Depth) of each object using depth rendering techniques [28], followed by volumetric fusion [10] to generate the complete TSDF grid $\mathbf{X}$. For the partial TSDF grid $\mathbf{c}$, we randomly select one of these views and fill out the TSDF voxel-grid using the camera pose and depth such that observed regions of space have a positive value while unobserved regions of space have the negative truncation value. The corresponding RGB image from this view is denoted as $\mathbf{I}$.
IV Network Architecture
Our diffusion model is fundamentally a U-Net [36] with an encoder, a middle block, and a decoder, which is connected to the encoder through skip connections. Both encoder and decoder contain ResNetBlocks. In the case of the encoder, these are accompanied by downsampling blocks, while in the decoder, they are accompanied by up-sampling convolution layers. The encoder, decoder, and middle block include two spatial transformers each. Images are encoded using CLIP encoder [30], and diffusion timesteps are encoded with a 2-layer MLP using positional encoding. The control branch shares the same architecture as the diffusion model up to the middle block and then uses 3D convolutional projection layers to feed partial shape’s $\mathbf{c}$ intermediate representation into the diffusion model in a spatially consistent manner. Tab. 6 shows the hyperparameters chosen in our experiments.
@@ IV Network Architecture
V More Qualitative Examples
[…]
Figure 10: Qualitative comparison of our method against state-of-the-art approaches on several samples from the ShapeNet dataset [6]. Compared to prior methods, our approach produces plausible, non-oversmoothed shapes while recovering fine details, such as the notch in the table (rightmost column).
[…]
Figure 11: Shape completion comparison with SDFusion [7] on samples from the ShapeNet dataset [6]. Our method produces more plausible and diverse results, whereas SDFusion [7] generates noisier outputs and tends to collapse to similar shapes.
Figure 12: Additional diverse predictions generated by our model. Note that it can produce completely separate modes, such as different backrest designs in the second row.
Figure 13: Examples of failure cases where our model does not accurately predict high-fidelity, realistic shapes. (a) The model predicts a container, although the ground-truth shape is a lamp. (b) The latent diffusion model unnecessarily generates a more complex shape than the ground truth. (c, d) For highly incomplete inputs, the model completes the shape as a different class—for instance, predicting a table and a sofa instead of the ground-truth lamp and cabinet.
@@ V More Qualitative Examples
TSDF Truncation threshold threshthresh $3.0$
@@
VQ-VAE latent feature size $D$ 33
@@
VQ-VAE 3D commitment loss weight $\lambda$ $0.5$
VQ-VAE 2D reconstruction loss weight $\gamma_{R}$ $0.4$
VQ-VAE 2D adversarial loss weight $\gamma_{A}$ $0.4$
@@
Diffusion timesteps at inference ([38]) $T_{inf}$ $100$
@@
Batch Size bsbs 88
Learning Rate - VQ-VAE $lr_{vqvae}$ $1e^{-4}$
Learning Rate - Diffusion Model $lr_{diff}$ $2.5e^{-5}$